# DESI EMISSION LINE GALAXIES: CLUSTERING DEPENDENCE ON STELLAR MASS AND [OII] LUMINOSITY

### v2

@@ 6 Conclusions
In this paper, we measure and model the clustering of ELGs from the DESI One-Percent Survey and its dependence on stellar mass and [OII] luminosity. In each of three redshift ranges ($1.00.8<z<$ 1.0, $1.21.0<z<$ 1.2, $1.61.2<z<$ 1.6), the projected 2PCFs and number densities of nine $\log L_{\rm[OII]}-\log M_{*}$ samples are measured. To fit the measurements of all samples at fixed redshift simultaneously, we apply a modified CCMD framework (Xu et al., 2018; Clontz et al., 2022) that describes the conditional $\log L_{\rm[OII]}-\log M_{*}$ distribution as a function of halo mass. We obtain constraints on physically-motivated model parameters, mean halo occupation functions, and derived quantities.
The model demonstrates flexibility in fitting the measurements across stellar mass and [OII] luminosity. Several features in the projected 2PCFs influence our fits. The 2PCF measurements plateau on small scales and have large uncertainties. The model responds with satellite fractions that are low or loosely constrained, making constraints on only the central galaxy parameters meaningful. On scales above $r_{p}\sim 10$ $h^{-1}$pc, a variety of spurious signal is observed in the 2PCFs, suggesting the presence of systematic errors in the imaging data and resulting selection of targets. This effect is most evident at low redshift, which leads to a $\chi^{2}$ value of 261.7 (93 degrees of freedom). Nonetheless, the measurements at intermediate and high redshift are well-fit. Finally, the 2PCFs show only weak dependence on galaxy property, which is largely reflected in the model constraints.
[…]
@@
Now that DESI has completed three years of observation, we are considering several modifications to our analysis for the larger data sample. Given the lack of strong property dependence found in this work, other bases beyond $\log L_{\rm[OII]}-\log M_{*}$ will be explored. In particular, directly observed properties such as magnitudes, colors, and morphology are of interest because they inform the DESI ELG target selection algorithm (Raichoor et al., 2023) and do not rely on modeling of the spectrophotometry. Joint sampling according to these properties will be examined in the upcoming DESI Year 3 sample to identify the basis with the most influence on the clustering measurements. Sample-dependent systematic weights should also be assigned to account for variations in the imaging quality. Such corrections may address the spurious signal that is observed in the projected 2PCF measurements at low redshift and stellar mass. The chosen basis may also inform changes to our model parameterization.
Regardless of basis, larger data releases by DESI will enable improved sample statistics and stricter quality selection. Utilizing the entirety of a larger sample will allow us to improve the small scale measurement and sample more finely in the galaxy properties. Alternatively, precision in the galaxy property estimates can be prioritized by using overlaps with deeper and more diverse photometry such as Data Release 3 of the Hyper Suprime-Cam Subaru Strategic Survey (HSC-SS; Aihara et al. 2022).
Finally, the model presented here relies on analytic descriptions of halo clustering (Jenkins et al., 2001; Navarro et al., 1996; Sheth et al., 2001; Tinker et al., 2005). With a larger data sample, we will turn to a more accurate and efficient simulation-based method to model the clustering measurement (Zheng & Guo, 2016; Xu et al., 2018).
Within the context of DESI, ELGs will constitute the largest fraction of the full five-year sample. To optimize cosmological constraints from the ELG sample, we should understand the physical properties of these tracers and the impact on their selection, clustering, and relation to the underlying matter distribution. Our constraints and future studies are catered specifically to DESI and can inform mock galaxy catalogs and galaxy evolution models for future analysis. Beyond DESI, continued investigation of the ELG-halo connection will remain critical as ELGs are leveraged to study the $z\gtrsim 1$ Universe in future spectroscopic surveys, such as those with Euclid (Laureijs et al., 2011) and the Nancy Grace Roman Space Telescope (Green et al., 2012).
@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science, Innovation and Universities of Spain (MICIU/AEI/10.13039/501100011033), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies.
@@
The COSMOS2020 data used here is based on observations collected at the European Southern Observatory under ESO programme ID 179.A-2005 and on data products produced by CALET and the Cambridge Astronomy Survey Unit on behalf of the UltraVISTA consortium.
@@
The work of T. H. and K. D. was supported in part by U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award No. DESC0009959. Z.Z. is supported by NSF grant AST-2007499.
@@
All data shown in figures are publicly available at https://doi.org/10.5281/zenodo.15467956.
@@ Appendix A Comparison with COSMOS2020
We show ELG stellar mass and [OII] luminosity estimates of DESI and COSMOS2020 in Figure 13. As a function of COSMOS2020 stellar mass, DESI stellar masses exhibit an offset of roughly -0.37 dex and a scatter of $\sigma_{\log M_{*}}\sim 0.5$ dex. With respect to DESI [OII] luminosities (applying the aperture correction in Section 2.3), the COSMOS2020 [OII] luminosities have an offset of -0.1 dex and a scatter of $\sigma_{\log L_{\rm[OII]}}\sim 0.13$ dex.
The large measurement scatter in stellar mass is interesting for several reasons. First, it exceeds the typical intrinsic scatter at fixed halo mass inferred from HOD modeling, $\sigma_{\log M_{*}}\sim 0.2$ (e.g., Wechsler & Tinker 2018). This measurement uncertainty is reflected in our model constraints, which capture the combined effect of intrinsic and measurement scatter. Second, such a large measurement scatter can flatten trends that might be expected in the $\log L_{\rm[OII]}-\log M_{*}$ plane. This effect is demonstrated in Figure 14, where DESI stellar masses (left) yield a flat distribution while COSMOS2020 stellar masses (right) yield a correlation. Whether a flat distribution should be expected is unclear. A slope may be observable via the star formation main sequence (e.g., Noeske et al. 2007; Daddi et al. 2007; Sparre et al. 2015), but [OII] luminosity is a limited proxy for star formation rate (Moustakas et al., 2006) and may be uncorrelated (Favole et al., 2020) or even anti-correlated (Favole et al., 2024) with stellar mass.
Figure 13: Comparison of ELG properties reported by DESI (via FastSpecFit) and COSMOS2020 (via EAZY). Left: DESI stellar masses have a notable offset and scatter with respect to COSMOS2020. DESI stellar masses have been converted to match COSMOS2020’s value of $h$. Both assume a Chabrier initial mass function (Chabrier, 2003). Right: DESI is now plotted on the horizontal axis, as the DESI [OII] luminosities are modeled from spectrophotometry and considered more accurate. COSMOS2020 [OII] luminosities have a small offset and scatter with respect to DESI.
Figure 14: DESI stellar masses (left) have a sufficiently large scatter to flatten the [OII] luminosity-stellar mass slope observed with COSMOS2020 stellar masses (right). This is also true at the higher redshifts considered in this work.
@@ Appendix B Aperture Correction
In this work, the measured flux is corrected using Equation (2) and $r$-band photometry from the DESI Legacy Imaging Surveys (Dey et al., 2019). An alternative correction can be derived by combining the equivalent widths (EW) of the [OII] doublet in the DESI spectra with the continuum flux predicted from the photometry,
@@
where $EW_{372X}$ is the equivalent width of the line at (rest-frame) 3,72X Å (measured from the spectrum) and $F_{3728}^{\rm continuum}$ is the modeled continuum flux around the [OII] doublet at (rest-frame) $\sim$3,728 Å (inferred from the photometry). All quantities are provided in the FastSpecFit value-added catalog (Moustakas et al., 2023).
Figure 15 compares the [OII] luminosities calculated with the $r$-based correction ($L_{r-\rm corrected}$) and the equivalent width-based correction ($L_{\rm EW-corrected}$). A few conclusions are drawn from Figure 15. In the left column, the one-to-one line at each redshift traces the median within one standard deviation and $\sim$0.1 dex. This suggests consistency in the central value between aperture corrections. However, the middle and right columns show that the distribution of $\log L_{\rm EW-corrected}$ at fixed $\log L_{r-\rm corrected}$ (right column) is more Gaussian than the reverse (middle column). This result suggests that the $r$-corrected [OII] luminosities are closer to truth. Further, the EW-based correction relies on a model to predict the continuum from only a few photometric measurements. This likely introduces more uncertainty than the assumed 1′′ FWHM seeing of the $r$-based correction. For these reasons, the $r$-based correction is preferred.
Figure 15: Comparison between aperture corrections. Each row contains one redshift group. Left column: Equivalent width corrected (“EW-corrected”) [OII] luminosity vs. $r$-band corrected (“$r$-corrected”) [OII] luminosity. Shaded squares and colorbar give the number of galaxies in each 2D bin. The one-to-one (black dashed) line traces the median (red solid) within one standard deviation and $\sim$0.1 dex, suggesting consistency in the central value. Middle column: Distribution of $r$-corrected $\log L_{\rm[OII]}$ at fixed EW-corrected $\log L_{\rm[OII]}$. Right column: Distribution of EW-corrected $\log L_{\rm[OII]}$ at fixed $r$-corrected $\log L_{\rm[OII]}$ (reverse of middle column).
@@ Appendix C Full Model Constraints
Figure 16: Constraints on all model parameters, for the low-redshift ($1.00.8<z<$ 1.0) fitting. In each corner plot, shaded boxes show the binned density of the model posterior. Solid contours enclose the central 68% and 95% of the 2D distributions. In each histogram, dashed vertical lines mark the median and central 68% interval.
Figure 17: Same as Figure 16 but for the $1.21.0<z<$ 1.2 intermediate-redshift fitting.
Figure 18: Same as Figure 16 but for the $1.61.2<z<$ 1.6 high-redshift fitting.

### v3

@@ 6 Conclusions
In this paper, we measure and model the clustering of ELGs from the DESI One-Percent Survey and its dependence on stellar mass and [OII] luminosity. In each of three redshift ranges ($0.8<z<1.0$, $1.0<z<1.2$, $1.2<z<1.6$), the projected 2PCFs and number densities of nine $\log L_{\rm[OII]}-\log M_{*}$ samples are measured. To fit the measurements of all samples at fixed redshift simultaneously, we apply a modified CCMD framework (Xu et al., 2018; Clontz et al., 2022) that describes the conditional $\log L_{\rm[OII]}-\log M_{*}$ distribution as a function of halo mass. We obtain constraints on physically motivated model parameters, mean halo occupation functions, and derived quantities.
The model demonstrates flexibility in fitting the measurements across stellar mass and [OII] luminosity. Several features in the projected 2PCFs influence our fits. The 2PCF measurements plateau on small scales and have large uncertainties. The model responds with satellite fractions that are low or loosely constrained, making constraints on only the central galaxy parameters meaningful. On scales above $r_{p}\sim 10$ $h^{-1}$Mpc, varying degrees of spurious signal are observed in the 2PCFs of different samples, suggesting the presence of systematic errors in the imaging data and resulting selection of targets. This effect is most evident at low redshift, which leads to a $\chi^{2}$ value of 261.7 (93 degrees of freedom). Nonetheless, the measurements at intermediate and high redshift are well fit. Finally, the 2PCFs show only weak dependence on galaxy properties, which is largely reflected in the model constraints.
We summarize our findings with three key takeaways. First, ELGs in the DESI One-Percent Survey reside in halos that span a narrow range in mass. In our work, that range is $\sim$$10^{12}$ $h^{-1}M_{\odot}-10^{12.7}$ $h^{-1}M_{\odot}$ and accompanied by low satellite fractions. When compared to other studies, the exact halo mass scale depends on the satellite fraction estimate, which largely explains discrepancies with, e.g., Prada et al. (2025). The existence of such a selection on halo mass is consistent with the occupation functions in other works (Gonzalez-Perez et al., 2020; Prada et al., 2025) and can be understood through the DESI target selection (Raichoor et al., 2023). The upper limit on $g$-band fiber magnitude removes faint ELGs that reside in low-mass halos, while cuts in grg-r and rzr-z colors exclude quiescent galaxies that dominate high-mass halos. In our model, this halo mass selection is reproduced by distorting the ELG central fraction ($A_{c}$) and median [OII] luminosity-halo mass relation ($\mu_{y}$) to remove galaxies from halos outside the narrow mass range.
@@
Now that DESI has completed 3 yr of observation, we are considering several modifications to our analysis for the larger data sample. Given the lack of strong property dependence found in this work, other bases beyond $\log L_{\rm[OII]}-\log M_{*}$ will be explored. In particular, directly observed properties such as magnitudes, colors, and morphology are of interest because they inform the DESI ELG target selection algorithm (Raichoor et al., 2023) and do not rely on modeling of the spectrophotometry. Joint sampling according to these properties will be examined in the upcoming DESI Year 3 sample to identify the basis with the most influence on the clustering measurements. Sample-dependent systematic weights should also be assigned to account for variations in the imaging quality. Such corrections may address the spurious signal that is observed in the projected 2PCF measurements at low redshift and stellar mass. The chosen basis may also inform changes to our model parameterization.
Regardless of basis, larger data releases by DESI will enable improved sample statistics and stricter quality selection. Utilizing the entirety of a larger sample will allow us to improve the small-scale measurement and sample more finely in the galaxy properties. Alternatively, precision in the galaxy property estimates can be prioritized by using overlaps with deeper and more diverse photometry such as Data Release 3 of the Hyper Suprime-Cam Subaru Strategic Survey (or HSC-SS; Aihara et al. 2022).
Finally, the model presented here relies on analytic descriptions of halo clustering (Navarro et al., 1996; Jenkins et al., 2001; Sheth et al., 2001; Tinker et al., 2005). With a larger data sample, we will turn to a more accurate and efficient simulation-based method to model the clustering measurement (Zheng & Guo, 2016; Xu et al., 2018).
Within the context of DESI, ELGs will constitute the largest fraction of the full 5 yr sample. To optimize cosmological constraints from the ELG sample, we should understand the physical properties of these tracers and the impact on their selection, clustering, and relation to the underlying matter distribution. Our constraints and future studies are catered specifically to DESI and can inform mock galaxy catalogs and galaxy evolution models for future analysis. Beyond DESI, continued investigation of the ELG-halo connection will remain critical as ELGs are leveraged to study the $z\gtrsim 1$ Universe in future spectroscopic surveys, such as those with Euclid (Laureijs et al., 2011) and the Nancy Grace Roman Space Telescope (Green et al., 2012).
@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility, under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences, under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science, Innovation and Universities of Spain (MICIU/AEI/10.13039/501100011033); and by the DESI Member Institutions (https://www.desi.lbl.gov/collaborating-institutions). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. National Science Foundation, the U.S. Department of Energy, or any of the listed funding agencies.
@@
The COSMOS2020 data used here are based on observations collected at the European Southern Observatory under ESO program ID 179.A-2005 and on data products produced by CALET and the Cambridge Astronomy Survey Unit on behalf of the UltraVISTA consortium.
@@
The work of T.H. and K.D. was supported in part by the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award No. DESC0009959. Z.Z. is supported by NSF grant AST-2007499.
@@
All data shown in figures are publicly available on Zenodo (doi:10.5281/zenodo.15467956).
@@ Appendix A Comparison with COSMOS2020
We show ELG stellar mass and [OII] luminosity estimates from DESI and COSMOS2020 in Figure 13 (left and middle panels, respectively). DESI stellar masses have been converted to match COSMOS2020’s value of $h$, and both assume a Chabrier initial mass function (Chabrier, 2003). As a function of COSMOS2020 stellar mass, DESI stellar masses exhibit an offset of roughly -0.37 dex and a scatter of $\sigma_{\log M_{*}}\sim 0.5$ dex. This DESI scatter exceeds the statistical uncertainty reported by COSMOS2020 ($\sim$0.04 dex on average) despite the COSMOS2020 estimates relying on photometric redshifts. The DESI stellar mass estimates are also skewed toward higher values at fixed COSMOS2020 stellar mass, and the extent of this tail is greatest at low stellar mass. With respect to DESI [OII] luminosities (applying the aperture correction in Section 2.3), the COSMOS2020 [OII] luminosities have an offset of -0.1 dex and a scatter of $\sigma_{\log L_{\rm[OII]}}\sim 0.13$ dex.
[…]
We also plot the COSMOS2020 star formation rate estimates against the DESI [OII] luminosities in Figure 13 (right panel) to inspect the validity of [OII] luminosity as a tracer for star formation rate. Here, star formation rate is positively correlated with DESI [OII] luminosity but has a scatter of $\sigma_{\log\rm SFR}\sim 0.23$ dex and skewness, suggesting that [OII] luminosity is a limited tracer of star formation. Other studies have also found limitations in [OII] luminosity as such a tracer (e.g., Moustakas et al. 2006).
The subtleties in the DESI ELG properties characterized above lead to consequences in our work, two of which we describe below.
First, the scatter in the DESI stellar mass estimates and limited validity of [OII] luminosity as an indicator of star formation rate suppress evidence of the star formation main sequence (e.g., Daddi et al. 2007; Noeske et al. 2007; Sparre et al. 2015) in the DESI $\log L_{\rm[OII]}-\log M_{*}$ plane. This effect is demonstrated in Figure 14, where an originally clear star formation main sequence (left panel) becomes shallower when using the noisier DESI stellar mass estimates (middle) and ultimately flattens in the DESI $\log L_{\rm[OII]}-\log M_{*}$ plane (right).
[…]
Second, the offset, scatter, and skewness of the DESI stellar mass distribution can affect our HOD modeling result. For example, the large stellar mass estimate uncertainty inflates our model constraints on the stellar mass scatter at fixed halo mass ($\sigma_{x}$ as defined in Section 4.1.1) beyond the intrinsic value of $\sigma_{\log M_{*}}\sim 0.2$ typically inferred from HOD modeling (e.g., Wechsler & Tinker 2018). Additionally, the skewness of the stellar mass estimates can alter the shape of the inferred SHMR.
@@ Appendix B Aperture Correction
In this work, the measured flux is corrected using Equation (2) and $r$-band photometry from the DESI Legacy Imaging Surveys (Dey et al., 2019). An alternative correction can be derived by combining the equivalent widths (EWs) of the [OII] doublet in the DESI spectra with the continuum flux predicted from the photometry:
@@
where $EW_{372\rm{X}}$ is the equivalent width of the line at (rest-frame) 372X Å (measured from the spectrum) and $F_{3728}^{\rm continuum}$ is the modeled continuum flux around the [OII] doublet at (rest-frame) $\sim$3728 Å (inferred from the photometry). All quantities are provided in the FastSpecFit value-added catalog (Moustakas et al., 2023).
Figure 15 compares the [OII] luminosities calculated with the $r$-based correction ($L_{r-\rm corrected}$) and the EW-based correction ($L_{\rm EW-corrected}$). A few conclusions are drawn from Figure 15. In the left column, the one-to-one line at each redshift traces the median within one standard deviation and $\sim$0.1 dex. This suggests consistency in the central value between aperture corrections. However, the middle and right columns show that the distribution of $\log L_{\rm EW-corrected}$ at fixed $\log L_{r-\rm corrected}$ (right column) is more Gaussian than the reverse (middle column). This result suggests that the $r$-corrected [OII] luminosities are closer to truth. Further, the EW-based correction relies on a model to predict the continuum from only a few photometric measurements. This likely introduces more uncertainty than the assumed 1′′ FWHM seeing of the $r$-based correction. For these reasons, the $r$-based correction is preferred.
Figure 15: Comparison between aperture corrections. Each row contains one redshift group. Left column: equivalent-width-corrected (“EW-corrected”) [OII] luminosity vs. $r$-band-corrected (“$r$-corrected”) [OII] luminosity. Shaded squares and color bar give the number of galaxies in each 2D bin. The one-to-one (black dashed) line traces the median (red solid) within 1 standard deviation and $\sim$0.1 dex, suggesting consistency in the central value. Middle column: distribution of $r$-corrected $\log L_{\rm[OII]}$ at fixed EW-corrected $\log L_{\rm[OII]}$. Right column: distribution of EW-corrected $\log L_{\rm[OII]}$ at fixed $r$-corrected $\log L_{\rm[OII]}$ (reverse of middle column).
@@ Appendix C Full Model Constraints
Figure 16: Constraints on all model parameters, for the low-redshift ($0.8<z<1.0$) fitting. In each corner plot, shaded boxes show the binned density of the model posterior. Solid contours enclose the central 68% and 95% of the 2D distributions. In each histogram, dashed vertical lines mark the median and central 68% interval.
Figure 17: Same as Figure 16 but for the $1.0<z<1.2$ intermediate-redshift fitting.
Figure 18: Same as Figure 16 but for the $1.2<z<1.6$ high-redshift fitting.